Lightbulbs of a certain type are advertised as having an average lifetime of 750 hours. The price of these bulbs is very favorable, so a potential customer has decided to go ahead with a purchase arrangement unless it can be conclusively demonstrated that the true average lifetime is smaller than what is advertised. A random sample of 50 bulbs was selected, and the sample mean was found to be 738.44 with a standard deviation of 38.20 and a standard error on the mean of 5.40. What is the p-value for this test (to 4 decimals)?
Answer #1

Given that,
population mean(u)=750
sample mean, x =738.44
standard deviation, s =38.2
number (n)=50
null, Ho: μ=750
alternate, H1: μ<750
level of significance, α = 0.05
from standard normal table,left tailed t α/2 =1.677
since our test is left-tailed
reject Ho, if to < -1.677
we use test statistic (t) = x-u/(s.d/sqrt(n))
to =738.44-750/(38.2/sqrt(50))
to =-2.14
| to | =2.14
critical value
the value of |t α| with n-1 = 49 d.f is 1.677
we got |to| =2.14 & | t α | =1.677
make decision
hence value of | to | > | t α| and here we reject Ho
p-value :left tail - Ha : ( p < -2.1398 ) = 0.01869
hence value of p0.05 > 0.01869,here we reject Ho
ANSWERS
---------------
null, Ho: μ=750
alternate, H1: μ<750
test statistic: -2.14
critical value: -1.677
decision: reject Ho
p-value: 0.01869 =0.0187
we have enough evidence to support the claim that the true average lifetime is smaller than what is advertised.